OpenStudy (anonymous):

b+c=21 si a=7, calculati axb+axc

OpenStudy (anonymous):

si a??

OpenStudy (anonymous):

ok, ab+ac=a(b+c)=21*7=147

OpenStudy (anonymous):

multumesc

OpenStudy (anonymous):

ok ab+ac=a(b+c)=21*7=147 now multiple in x so it will be 147x in other word it is a function of x